### v1

# LoL-PIM: Long-Context LLM Decoding with Scalable DRAM-PIM System

Abstract.
The expansion of large language models (LLMs) with hundreds of billions of parameters presents significant challenges to computational resources, particularly data movement and memory bandwidth. Long-context LLMs, which process sequences of tens of thousands of tokens, further increase the demand on the memory system as the complexity in attention layers and key-value cache sizes is proportional to the context length. Processing-in-Memory (PIM) maximizes memory bandwidth by moving compute to the data and can address the memory bandwidth challenges; however, PIM is not necessarily scalable to accelerate long-context LLM because of limited per-module memory capacity and the inflexibility of fixed-functional unit PIM architecture and static memory management. In this work, we propose LoL-PIM– a multi-node PIM architecture that accelerates long context LLM through hardware-software co-design. In particular, we propose how pipeline parallelism can be exploited across a multi-PIM module while a direct PIM access (DPA) controller (or DMA for PIM) is proposed that enables dynamic PIM memory management and results in efficient PIM utilization across a diverse range of context length. We developed an MLIR-based compiler for LoL-PIM– a commercial PIM-based compiler where the software modifications were implemented and evaluated, while the hardware changes were modeled in the simulator. Our evaluations demonstrate that LoL-PIM significantly improves throughput and reduces latency for long-context LLM inference, outperforming both multi-GPU and GPU-PIM systems (up to 8.54$\times$ and 16.0$\times$ speedup , respectively), thereby enabling more efficient deployment of LLMs in real-world applications.
*Equal contribution.
copyright: nonecopyright: acmlicensedjournalyear: 2025
1. Introduction
The rapid advancement of large language models (LLMs) has transformed fields ranging from natural language processing to intelligent agents by enabling the generation of contextually relevant responses across diverse applications  (OpenAI et al., 2024; Team et al., 2023; Touvron et al., 2023; Claude, 2024; Jiang et al., 2023). In particular, long-context LLMs, capable of maintaining coherence across tens of thousands of tokens, have significantly enhanced contextual relevance in various tasks. For instance, long document summarization  (Zhong et al., 2021) generates cohesive summaries from dispersed information across different sections of extensive text, while repository-level code analysis  (Liu et al., 2023) extends programming assistants’ capabilities to analyze entire codebases comprising thousands of lines. Furthermore, chain-of-thought (CoT) reasoning has recently improved answer quality by leveraging multi-step contextual reasoning  (Wei et al., 2022; OpenAI, 2024). These advancements rely on inference-time processing of extended contexts, with recent LLMs supporting long-context handling capabilities ranging from 16K to 200K tokens  (Bai et al., 2023; MosaicML, 2023; Claude, 2024; OpenAI, 2024).
There have been many recent works on accelerating LLM inference, but most have focused on LLM inferences with shorter context (i.e., up to 4 or 8k) (Hong et al., 2023b; Zeng et al., 2024; Wu et al., 2024b; Aminabadi et al., 2022). As the context length increases, both the memory capacity and the memory bandwidth become a greater bottleneck during LLM inference. In particular, the memory capacity for the key-value cache(KV-cache), which is needed during Attention computation, increases linearly with the input context length. Since Attention relies on general matrix-vector multiplication (GEMV), the compute density (operations per byte) of LLM decoding is reduced and decoding becomes memory bandwidth bound. To provide sufficient memory capacity and bandwidth, multi-GPUs are commonly used but results in underutilization of GPU compute.
To address the memory bandwidth challenges, DRAM-based Processing-in-Memory (PIM) technology offers a promising solution by embedding computational capabilities directly within the memory (Asifuzzaman et al., 2023; Mutlu et al., 2020; Lee et al., 2021b; Kwon et al., 2023b; Lavenier et al., 2016; Choi et al., 2023; Park et al., 2024; Heo et al., 2024). Fixed-functional unit PIM systems, such as AiMX (Lee et al., 2022; Kwon et al., 2023b; He et al., 2020; Kwon et al., 2022, 2023a), are particularly well-suited for accelerating GEMV in LLM inference by exploiting the high internal memory bandwidth within DRAM modules. While PIM can accelerate GEMV, leveraging PIM for long-context LLM decoding presents new challenges, including the following.
  1. (1)
    Memory capacity: Individual PIM chip memory capacity is limited to a few GBs and is problematic for large KV-caches that require up to hundreds of GBs for long context.
  2. (2)
    Static PIM memory management: Long-context LLM have a diverse set of input context lengths and a fixed (or static) memory management results in inefficient usage of PIM memory and reduces overall throughput.
  3. (3)
    I/O buffer bottleneck: Non-square, aspect ratio of KV-cache, caused by long context inputs, creates a new PIM bottleneck in the I/O buffers.
Recent work (Park et al., 2024; Heo et al., 2024; Seo et al., 2024) exploit PIM and heterogeneous architectures (e.g., GPU-PIM or NPU-PIM) to accelerate LLM inference; however, they focus on short-context LLM (under 2k tokens) and do not address the challenges of long-context LLM. In particular, prior work often have assumed fixed context length when serving multiple requests; however, making the same assumption for long-context LLM results in poor utilization of the PIM since PIM memory would need to be allocated for the maximum context length, which can be an order of magnitude higher than the average context length. To overcome the challenges of accelerating long-context LLM, this work proposes LoL-PIM, a novel scalable multi-node PIM system architecture for long-context LLM that supports up to 32K tokens. Each node in LoL-PIM consists of multiple PIM modules to provide sufficient memory capacity (and bandwidth) to address the challenges of long-context LLM acceleration. PIM-aware partitioning of the KV-cache and weight matrix is proposed to enable sequence parallelism within the PIM modules to maximize bandwidth utilization and improve decoding throughput (tokens/sec).
Modern PIM (Lee et al., 2021b; Kwon et al., 2023a) architectures with fixed-functional units are often restricted to static PIM memory management because of limited programmability. This is problematic for long-context LLM since each request must be allocated based on the maximum context length. To address this limitation, LoL-PIM proposes dynamic memory management through Direct PIM Access (DPA) controller, or DMA for PIM, to support diverse KV-cache sizes and enable “lazy” PIM memory allocation. This results in efficient usage of PIM memory capacity for long-context LLM while also supporting larger batch sizes to improve overall throughput. Both PIM-aware data partitioning and dynamic memory management maximize internal memory bandwidth and ensure efficient utilization of PIM memory but create a new bottleneck in the PIM I/O. Thus, LoL-PIM includes a novel PIM I/O-aware buffering to mitigate the I/O bottleneck in long-context LLM. In particular, ping-pong I/O buffer is proposed which double buffers PIM input and output to overlap I/O data movement with PIM computation.
Figure 1. Decoding Computation for Long-Context LLM.
[…]
We evaluate LoL-PIM’s scalable decoding performance on popular long-context LLMs (7B to 72B parameters) using long-context tasks (token length ranging from 2K to 32K) from a representative benchmark, LongBench (Bai et al., 2024). We evaluate LoL-PIM as both PIM-only system or a heterogeneous system with a GPU. Our evaluations show
LoL-PIM outperforms the baseline GPU and the commercial AiMX in token generation throughput by 8.54$\times$ and 4.74$\times$.
2. Background
Figure 2. Multi-node PIM system and parallel LLM execution. (a) A multi-node PIM system organization. (b) A tensor parallel (TP) and pipeline parallel (PP) LLM execution. (c) The execution flow of PIM commands for LLM inference.
[…]
2.1. Long-context LLM Decoding
Long-context LLM Inference. Long-context LLMs excel at maintaining coherence over tens of thousands of tokens, significantly enhancing contextual relevance across various tasks – e.g., long document summarization (Zhong et al., 2021) generates cohesive summaries from information spread across different sections of extensive text, and repository-level code analysis (Liu et al., 2023) extends programming assistants’ capabilities to analyze entire codebases comprising thousands of lines. Chain-of-thought (CoT) reasoning also improves answer quality by leveraging multi-step contextual reasoning (Wei et al., 2022; OpenAI, 2024). These advancements rely on inference-time processing of extended contexts, with modern LLMs supporting token lengths ranging from 16K to 200K (Bai et al., 2023; MosaicML, 2023; Claude, 2024; OpenAI, 2024). Table 1 summarizes publicly available long-context LLMs, including the number of layers ($n_{l}$), attention heads ($n_{h}$), and the feature dimension of each head ($d_{h}$), as well as structural variations like group query attention (GQA)(Ainslie et al., 2023) and SwiGLU(Shazeer, 2020). (Implementation details are discussed in Sec.7). These models are trained on long-context datasets, enabling them to perform tasks requiring extended input contexts processing. Table 2 highlights the characteristics of popular long-context benchmarks, which exhibit significant variability in context lengths across requests.
Model $n_{l}$ $n_{h}$ $d_{h}$ SwiGLU GQA Reference CTL LLM-1.8B 24 16 64 $\times$ Qwen1.5-1.8B (Bai et al., 2023) 32K LLM-7B 32 32 128 $\times$ (ReLU) $\times$ MPT-7B-story (MosaicML, 2023) 65K $\times$ Qwen1.5-7B (Bai et al., 2023) 32K Llama3.1-8B (Dubey et al., 2024) 128K LLM-14B 40 40 128 $\times$ Qwen1.5-14B (Bai et al., 2023) 32K LongChat-13B (Li* et al., 2023) 16K LLM-72B 80 64 128 $\times$ Qwen1.5-72B (Bai et al., 2023) 32K Llama3.1-70B (Dubey et al., 2024) 128K
Table 1. Specifications and context length(CTL) limits of various Long-Context LLMs.
Task QMSum HotpotQA Musique mean $13966$ $13465$ $16362$ std $6182$ $3921$ $1651$ max $30456$ $17674$ $17917$ min $2651$ $1917$ $6820$
Table 2. Statistics of input context length with tokenizer of Qwen Model (LongBench (Bai et al., 2024)).
Decoding Computation. Long-context large language models (LLMs) are constructed based on the Transformer decoder (Vaswani et al., 2017). As illustrated in Fig. 1, each decoder of $n_{l}$ layers comprises Multi-Head Attention (MHA) and Feed-Forward Network (FFN). Each head $h$ of MHA starts with the Query/Key/Value generation (QKV-Gen) that generates vectors $\{Q,K,V\}\in\mathbb{R}^{d_{h}}$, where $K$ and $V$ are stacked to KV-cache $\in\mathbb{R}^{t\times d_{h}}$ for $t$ tokens. It is followed by the score generation ($QK^{T}$), which is then scaled and softmaxed ($S=\text{softmax}(QK^{T}\sqrt{d_{h}})$) to generate the scaled-dot-product attention ($SV$). Then, $SV$ concatenated across $n_{h}$ heads is projected (Proj) to become the MHA output. FFN consists of two consecutive fully-connected (FC) layers, FFN1 and FFN2, with an activation function in between for non-linear transformation. In sum, the Transformer decoder computation can be categorized into two types: FC-layers (QKV-Gen, Proj, FFN1, FFN2) that involve weight matrix, and attention-layers ($QK^{T}$ and $SV$) that involve KV-cache.
2.2. Multi-node Domain-specific PIM System
DRAM-based Domain-specific PIM. DRAM-based Processing-in-Memory (PIM) technology addresses memory-bound challenges by integrating computational capabilities directly within the memory subsystem (Asifuzzaman et al., 2023; Mutlu et al., 2020; Lee et al., 2021b; Kwon et al., 2023b; Lavenier et al., 2016; Choi et al., 2023; Park et al., 2024; Heo et al., 2024). Fixed functional unit PIM systems are optimized for accelerating LLM inference through GEMV operations. These PIM feature processing units adjacent to DRAM banks for parallel dot-product computations, pre-generated PIM commands for simplified control, and limited conditional logic to minimize area overhead  (Lee et al., 2022; Kwon et al., 2023b; He et al., 2020; Kwon et al., 2022, 2023a). For example, AiMX (Lee et al., 2022; Kwon et al., 2023b; He et al., 2020; Kwon et al., 2022, 2023a) includes a vector multiply-accumulate unit in each DRAM bank for 16-element dot-product computations, a shared 2KB global buffer for input data, and a minimal pair of 2-byte output registers per processing unit to transfer results off-chip for module-level processing (e.g., Softmax) using Extra Processing Unit (Fig. 2(a)). Additionally, all channels share a broadcasting input data path to minimize resource usage.
PIM Cmd Description Arguments Ref. Cmd (Kwon et al., 2023b) WR-INP Copy Input from GPR to GB GPR index WRGB DOT-PROD Dot-Product on a DRAM row Row/Col index MACABK RD-OUT Copy Output from OutReg to GPR GPR index RDMAC
Table 3. PIM commands for mapping LLM inference.
Multi-node PIM System. Fig. 2 (a) illustrates the LLM inference on a multi-node system. The system consists of a host CPU connected with multiple nodes, each consisting of a compute-intensive xPU (i.e., GPU or NPU) connected with multiple PIM modules that provide extensive internal memory bandwidth. As motivated by prior works (Park et al., 2024; Heo et al., 2024; Seo et al., 2024), xPU takes responsibility for the compute-intensive context summarization phase (prefill), while PIM modules handle the memory-bandwidth/capacity-demanding token generation phase (decoding). PIM Controller supports program execution for LLM inference compiled for model-parallel serving. Tensor-parallel (TP) processing (Shoeybi et al., 2019) performs head-wise partitionining of MHA and FFN computations to minimize synchronization, while pipelined parallelism (PP) (Huang et al., 2019; Ma et al., 2024) groups the layers and assigns a node to each layer-group for batch-wise parallel processing. As an example, Fig. 2(b) illustrates a parallel execution of four inference requests in a batch are partitioned over four PIM modules (i.e., $(TP,PP)=(2,2)$), being dynamically managed as the request (= Request1) with the end-of-sequence (EOS) token is replaced with a new request (= Request5).
Execution Flow. Executing LLM inference on PIM hardware involves generating PIM commands (Table 3) to manage tasks such as matrix multiplications and attention mechanisms. Fig. 2(c) illustrates the execution flow, where the PIM module interacts with DRAM to store KV caches and directly compute dot-products. These PIM commands incorporate key configuration details such as layers, attention heads, and batch size. However, since the number of tokens during inference is unknown at compile time, PIM commands must be pre-generated for all potential token ranges up to the maximum size, and memory must be pre-allocated accordingly.
3. Motivation
3.1. Characteristics of Long-context LLM
Long-context LLM inference presents significant challenges due to the memory-intensive demands of Attention layers, particularly during token generation for batches of tens to hundreds of requests in the decoding phase. To analyze workload characteristics, we evaluate a long-context LLM (Qwen1.5-7B) on an A100-80GB GPU with input context lengths ($L_{in}$) ranging from 4K to 32K tokens. Weak scaling is used – thus, as the number of GPUs is scaled to ensure sufficient memory capacity, the batch size is maximized based on total GPU memory. The key observations are as follows:
  • KV-cache Memory Dominance: The KV-cache consumes a far greater portion of memory compared to weight parameters, as its size is proportional to the context length. Each token generation requires a prefilled KV-cache for Attention computations, leading to significant memory usage as the context length increases (Fig. 3(a)).
  • Attention Bottleneck: The growing computational demands of Attention become the major bottleneck at longer context lengths. Attention computations rely on general matrix-vector multiplication (GEMV), which reduces compute density (operations per byte) and makes decoding heavily reliant on memory bandwidth. While batch processing enhances throughput for fully connected (FC) layers, which use general matrix-matrix multiplication (GEMM), it exacerbates bandwidth limitations for Attention layers (Fig. 3(b)).
  • GPU Scalability: A single GPU lacks the memory capacity and bandwidth required to support long-context LLM demands (Agrawal et al., 2024; Lin et al., 2024; Wang et al., 2024; Wu et al., 2024a; Pawar et al., 2024). Multi-GPU can provide memory capacity and bandwidth but it results in underutilized computational resources and increased inference costs (tokens per dollar) that scale linearly with the number of GPU nodes (Fig. 3(c)).
Figure 3. Characteristics of LLM workloads with scaling context length. The x-axis represents the number of GPUs and input context lengths($L_{in}$). (a) Required memory capacity. (b) Floating-point operations and Ops/Byte. (c) GPU utilization and cost, normalized across configurations.
[…]
3.2. Challenges of PIM Scaling
In this subsection, we highlight the challenges of scaling fixed-functional unit PIM architecture for long-context LLM decoding. First, bank-parallel dot-product efficiency of PIM relies on matrix partitioning, but KV-caches may exceed the capacity of a single PIM module. Increasing the number of PIM modules to accommodate longer contexts raises $TP$, distorting the partitioned subsection’s aspect ratio and incurring frequent input/output transfers that lead to I/O transfer overhead. Fig. 4(a) illustrates the scaling of context length and node count. In GPU systems, throughput suffers from low compute utilization caused by memory bottlenecks. While PIM improves throughput in LLM decoding, its effectiveness diminishes as $TP$-only parallelization limits scaling. This work identifies inefficiencies in $PP$ from existing partitioning approaches and introduces a new strategy that enhances bank-parallel computation by enabling $PP$-friendly partitioning (Sec. 4).
Second, varying context lengths in decoding requests require dynamic PIM memory management, but fixed-functional PIM units have limited conditional control to maximize compute density. Commercial PIM with fixed-functional units are managed by pre-generated commands with embedded operand addresses, forcing memory allocation based on the maximum context length regardless of actual context length, leading to underutilized memory and constrained batch sizes. Fig. 4(b) demonstrates slow batch size growth in PIM systems. To address this, we propose a dynamic memory management method using lightweight dynamic control commands for lazy memory allocation, significantly increasing average batch size. As shown in Fig. 4(b), the proposed method achieves near-ideal batch sizes (Sec. 5).
Third, the KV-cache’s structure suffers from distortion due to low $d_{h}$, while the aspect ratio of weight parameters becomes increasingly biased as more PIM modules partition them. This biased matrix aspect ratio limits input and output reuse. Fixed-functional PIMs exacerbate this by constraining data transfer paths, resulting in significant I/O overhead. To mitigate this, we propose an I/O-aware buffering mechanism that overlaps data transfer with PIM computation. The resource overhead is minimal, as channels or banks share the buffers. Fig. 4(c) shows a substantial reduction in I/O overhead across varying context lengths (Sec. 6).
Figure 4. Performance evaluation of the proposed LoL-PIM compared to baseline-PIM and GPU implementations. The x-axis represents the system’s memory capacity, and the workload (context length: $L_{in}$) assigned to each case increases proportionally with the system’s capacity, like Fig.3 (Capacity = 80GB × N, Context length = 4K × N).
[…]
Figure 5. (a) PIM-aware partitioning. (b) Challenges for PIM-based pipeline parallelization. (c) Intra-module token-parallel partitioning.
[…]
4. Scalable PIM Architecture
4.1. PIM-aware Partitioning
LLM decoding workloads rely on weight parameter matrices and the KV-cache. Workload partitioning ensures that the dimensions of the matrix are distributed across PIM modules, with each subsection fitting within the memory capacity of the module: $N_{ch}\cdot N_{bank}\times D_{row}\cdot N_{row}$. When partitioning LLM workloads across multiple PIM modules, two primary parallelization methods are used: Pipeline Parallelization (PP) (Huang et al., 2019; Ma et al., 2024), which groups LLM layers into $PP$ pipeline stages, and Tensor Parallelization (TP) (Shoeybi et al., 2019), which divides matrix dimensions within each group by $TP$. While both approaches are widely applied in LLM inference systems, TP is more commonly preferred in multi-node PIM systems. A widely used strategy in prior works, called head-first allocation (HFA) (Heo et al., 2024; Seo et al., 2024; Park et al., 2024), is summarized below (Fig. 5(a)):
  • Partitioning Weight Parameters. In FC layers, weight parameters are typically partitioned by $TP$: $d_{out}$ for $QKV$ and $FFN1$, and $d_{in}$ for $Proj$ and $FFN2$, as per Tensor Parallelization (Shoeybi et al., 2019).
  • Partitioning KV-cache. In Attention computation, the dimensions $n_{h}$ and $B$ are independent and parallelized by $TP$, ensuring each request’s per-head KV-cache resides within a single channel (Heo et al., 2024; Park et al., 2024).
Increasing the number of PIM modules to support longer contexts raises $TP$, but this approach introduces inefficiencies (Fig. 5(a)): 1) In FC layers, excessive $TP$ distorts the partitioned subsection’s aspect ratio, causing frequent input/output transfers and resulting in I/O overhead. 2) In Attention layers, increasing $TP$ reduces channel occupancy, leading to bank underutilization.
4.2. Challenges for PIM-based Pipeline Parallelization
As the number of PIM modules increases, $PP$ becomes essential to address inefficiencies in bandwidth utilization caused by overly aggressive $TP$. However, using $PP$ with the HFA partitioning method presents a challenging performance trade-off. As shown in Fig. 5(b), each PIM module processes a micro-batch ($B_{\mu}$) of the KV-cache during Attention computation. For a given batch size $B$, determined by the memory capacity for static max-context-length allocation, the number of batch processing steps is $B/B_{\mu}$.
Pipeline bubbles, caused by host-PIM synchronization, grow as $B_{\mu}$ increases, leading to significant overhead. Conversely, smaller $B_{\mu}$ values result in fewer banks being used for parallel processing, causing underutilization. This trade-off has made $PP$ less favorable in multi-PIM parallelization, as noted in (Heo et al., 2024).
Figure 6. Direct PIM Access (DPA) controller. (a) DPA-encoded PIM commands. (b) DPA decoding example. (c) On-module dispatcher for supporting DPA. (d) Lazy memory allocation.
[…]
4.3. Intra-module Token-parallel Partitioning
Achieving balanced channel utilization requires minimizing batch dimension partitioning during KV-cache allocation. To address this, we propose a new strategy called intra-module token-parallel partitioning (ITPP). As shown in Fig. 5(c), this approach prioritizes the partitioning of output dimensions over the batch dimension, allocating the KV-cache based on their respective output dimensions: the token dimension for the Key cache and the head dimension for the Value cache.
ITPP differs from prior strategies in two key ways. First, it allocates the token dimension for bank-level parallelization, leveraging the abundant token dimension in long-context LLM workloads to avoid bank underutilization. Unlike HFA, which over-partitions the batch dimension and suffers from PP trade-offs, ITPP ensures more efficient utilization. Second, ITPP aggregates token-parallel outputs of $QK^{T}$ within a module, enabling the Extra Processing Unit (EPU) in the PIM Controller Hub to compute the Softmax aggregation. This allows for head-wise pipelined execution of the Softmax and subsequent $SV$ computation without incurring additional overhead. In contrast, Flash-Decoding (Dao et al., 2023; Hong et al., 2023a) involves numerically unstable partial Softmax computation (Golden et al., 2024) and costly inter-module communication. By combining the parallel computation advantages of Flash-Decoding for $QK^{T}$ and the effective $SV$ handling of HFA, ITPP delivers the best of both approaches. In summary, the benefits of the ITPP strategy include the following:
  • Balancing KV-cache Distribution. HFA causes an imbalanced KV-cache distribution for storing long-context KV cache, leaving some channels within a PIM module idle while overloading others, which underutilizes bandwidth. In contrast, ITPP uses token-parallel partitioning to evenly distribute workloads, mitigate the effects of context-length variability, and minimize channel imbalances.
  • Consistent Channel Utilization. By avoiding excessive partitioning of the batch dimension via token-paralle, ITPP ensures consistent utilization of all channels, even in cases with small batch sizes. By integrating output dimension prioritization and token dimension partitioning, LoL-PIM achieves efficient intra-module workload distribution without introducing additional overhead to attention operations.
  • Improved Bandwidth Utilization for FC-Layers. ITPP enhances bandwidth utilization for FC layers by enabling pipeline parallelism (PP) in combination with tensor parallelism (TP), reducing excessive $TP$ workload division.
5. Dynamic PIM Memory Management
Managing memory in long-context LLM decoding is challenging due to variability in context lengths. Although GPU-based approaches (Druschel et al., 2023) have implemented dynamic memory allocation for such variability, PIM’s static control capabilities limit its adaptability. To overcome this, we introduce the Direct PIM Access (DPA) controller, or DMA for PIM, which supports diverse KV-cache sizes and enables “lazy” memory allocation. DPA employs a lightweight set of commands and a command dispatching logic to facilitate dynamic memory management, significantly improving throughput in long-context LLM decoding.
5.1. Managing KV-cache in PIM
PIM accelerates Attention computations by managing the KV-cache within its memory modules. Recent studies (Park et al., 2024; Heo et al., 2024; Seo et al., 2024) have used static memory allocation for the KV-cache, based on the maximum token length, and pre-generated PIM commands with fixed physical addresses (e.g., DOT-PROD(row,col) in Table 3). However, PIM lacks flow control instructions, requiring the compiler to pre-generate commands for all possible token lengths and select the appropriate one at runtime. This approach results in inefficient static memory allocation and token-length-agnostic command generation. Pre-generated commands with fixed operand indices cannot adapt to varying KV-cache sizes encountered during LLM decoding. To address these issues, we make two key observations:
DPA Cmd. Arguments Description Dyn-Loop Loop-Bound (LB) Num. PIM CMD repitition. Loop-Entry (LE) Num. PIM commands that will be repeated Dyn-Modi Target Target bit-field to adjust (eg. Row address) Coefficient Value to modify target field
Table 4. Direct PIM Access (DPA) Commands.
[…]
  • Repetitive Compute Patterns: The computation sequence for Attention operations (e.g., $QK^{T}$, $SV$) follows a repetitive pattern across layers and tokens. As shown in Fig. 6(a), the sequence consists of identical PIM commands, differing only in the number of repetitions (e.g., repeated DOT-PROD and RD-OUT commands as token length increases). This suggests the possibility of dynamically unrolling PIM commands by encoding the repetition count.
  • Token-length-dependent Operands: Operand indices for PIM commands are determined solely by token length. For example, DOT-PROD operand indices (row) are calculated as $row=T_{cur}/256$, derived from matrix partitioning. The Va2Pa table provides a one-to-one mapping between virtual addresses (Va) and physical addresses (Pa), enabling dynamic identification of the KV-cache data address. Virtual addresses (Va) are assigned logically and continuously, starting from 0 and incrementing by 1 for each data entry. For instance, in Request 1, Va=0 maps to Pa=22, while in Request 2, Va=0 maps to Pa=33. By using the physical address from this mapping, the KV-cache data address (e.g., row=22 or row=33) can be efficiently identified on-the-fly.
Building on these insights, we propose a DPA commands to encode both repetition counts and operand modifications into a compact PIM command stack. This enables dynamic adjustment of PIM physical addresses to allocate and access KV-cache with varying context lengths during runtime. To implement this, we introduce a new PIM command set that includes repetition counts and operand updates. Additionally, we enhance the PIM module’s instruction dispatching logic to decode these commands and dynamically generate sequences at runtime. Together, these innovations provide a scalable and efficient solution for managing KV-cache in PIM environments, effectively addressing challenges posed by dynamic context lengths.
5.2. Direct PIM Access (DPA) Commands
Table 4 outlines the DPA commands designed to enhance the flexibility of PIM commands. These commands encapsulate repetitive token operations, enabling runtime execution without requiring re-compilation. The Dyn-Loop command encodes a loop structure, with Loop-Bound (LB) specifying the number of repetitions and Loop-Entry (LE) defining the number of commands to repeat. These parameters are determined dynamically based on the token length of each request. The Dyn-Modi command adjusts target operand fields (e.g., Target=row for DOT-PROD) of the subsequent PIM command using a coefficient as the stride (e.g., Coef=1 for row-wise traversal of DOT-PROD). These modified operands function as virtual addresses, initialized at zero and incremented by the stride during runtime dispatch. The request ID and virtual address are then used to retrieve the corresponding physical address from the virtual-to-physical memory table. For example, as shown in Fig. 6(c), for Request-2 with $T_{cur}=300$, LB=2 and LE=4. The first DOT-PROD command maps to the physical address (row=33), and the second maps to (row=34). This demonstrates how PIM commands dynamically determine operand physical addresses during runtime execution.
5.3. On-module PIM Command Dispatcher
To support DPA, we enhance the Instruction Sequencer within the PIM Controller (Fig. 2) to enable runtime dispatching of PIM commands. The on-module dispatcher manages dynamic command generation and memory allocation locally within the PIM module. Fig. 6(c) illustrates its components and operation, which include a Configuration Buffer, a Virtual-to-Physical Address (Va2Pa) Table, a Command Buffer, and a Decoding Unit. The total buffer size required for the dispatcher is minimal, less than 200KB(Table 5, which is significantly smaller than the 512KB GPR capacity in typical PIM Control Hubs (Yongkee Kwon et al., 2024), ensuring efficient integration. The Configuration Buffer stores decoding information such as the total number of layers, the current layer ID, request ID, and the token index ($T_{cur}$). The Va2Pa table and the Command Buffer are initialized with the virtual-to-physical address mapping of the KV-cache and the DPA-encoded PIM command stack, respectively. The host updates the Configuration Buffer each iteration with current decoding information, such as $T_{cur}$. Upon completing a request, the allocated memory chunk is released by updating the Va2Pa table, while the new request ID and associated $T_{cur}$ are received. The Decoding Unit uses the Configuration Buffer, Command Buffer, and Va2Pa table to dispatch the corresponding PIM command sequence.
5.4. Use Case: Lazy Memory Allocation
Integrating the proposed DPA and on-module dispatcher enables lazy memory allocation, which dynamically adjusts PIM memory capacity based on requests’ context lengths. The KV-cache increases with each iteration and eventually exceeds the pre-allocated chunk capacity. At this point, the Va2Pa table is updated to allocate a new chunk to accommodate the additional KV-cache. As shown in Fig. 6(d), when Request-1 requires more capacity, the host allocates a new chunk and updates the Va2Pa table. Notably, the new chunk does not need to be adjacent to the previous one, allowing on-demand allocation in non-contiguous memory regions and improving memory utilization.
Lazy memory allocation also increases the batch size for pipeline parallelization. The host determines the batch size for each iteration of pipeline processing. Under static memory allocation, requests must be assigned memory sufficient for the maximum context length, which limits the number of requests that can be processed simultaneously, even when current requests have smaller context lengths. In contrast, lazy memory allocation uses a smaller number of non-contiguous chunks to handle requests with small KV-cache sizes, significantly boosting average batch size. For example, the average batch size increases by up to 380% compared to static allocation, approaching near-optimal levels (Fig. 4). As discussed in Sec. 4, a larger effective batch size improves PIM’s capacity utilization, reduces pipeline bubbles, and enhances throughput.
Figure 7. (a) Comparison of latency breakdowns for each LLM operation with and without I/O-aware buffering. (b) Illustration of the implementation details for I/O-aware buffering, including data flow and the ping-pong bit mechanism.
[…]
6. PIM I/O Bottleneck
I/O Bottleneck Analysis. While our mapping strategy improves LoL-PIM’s efficiency for LLM inference, it doesn’t address the I/O data transfer overhead inherent to it. LoL-PIM’s buffers are significantly smaller than those in recent DRAM-PIM architectures (Heo et al., 2024; Park et al., 2024), limiting input and output data reuse. Fig. 7(a) reveals that data transfer latency for Out-Reg (DT-Out) and GB (DT-GB) accounts for over 50% of total latency in $QK^{T}$ and $SV$ operations. This highlights the need for architectural improvements in LoL-PIM to reduce input and output data transfer overhead.
I/O-aware buffering for Overcoming I/O Bottleneck. We propose I/O-aware buffering, strategy for LoL-PIM to address data transfer bottlenecks. It uses two buffer sets: one for current dot-product operations and another for preparing for the next, hiding transfer cycles within the computation. Fig. 7(b) illustrates I/O-aware buffering applied to input and output transfers via Global Buffer (GB) and Out-Reg. This approach maintains LoL-PIM’s main computation datapath with minimal hardware changes: extended control logic (a pingpong-bit for alternating buffers), additional muxes for buffer selection, and an extra GB shared across banks. I/O-aware buffering leverages existing Out-Reg sets, suiting PIM’s constrained environment.
Evaluation. To evaluate the proposed method, we modified LoL-PIM’s control path in Ramulator (Kim et al., 2016), implementing a dual command queue and pingpong-bit mechanism. This allows intermittent execution blocking and bit reversal for subsequent commands. The MAC units’ datapath remains unchanged, preserving critical path elements. The main overhead comes from the controller’s pingpong-bit-based command execution and an extra GB. Area estimation using CACTI and Synopsys Design Compiler with SAED 32nm technology shows only a 1.7% increase for a 256-MAC unit. Fig. 7(a) demonstrates I/O-aware buffering’s significant latency reductions: 40% in $QK^{T}$ (overlapping DT-Out with DOT-PROD(MAC) latency), 44% in SV (overlapping DT-GB with DOT-PROD latency), 29% in $FFN^{1}$ and 28% in $FFN^{2}$.
Figure 8. Overall compilation flow of LoL-PIM.
[…]
7. LoL-PIM System Architecture
We implement LoL-PIM as a software-hardware codesign to enable the proposed ideas of PIM-aware partitioning, dynamic memory management, and I/O aware buffering to enhance PIM’s capability for long-context LLM decoding. We extend the existing machine learning compiler framework to integrate PIM-specific compilation flow and code generation for diverse variants of Transformer decoding workloads. Also, we augment AiMX’s microarchitecture to enable on-module command dispatching and I/O aware buffering.
7.1. Multi-PIM Compiler and Runtime
We extend a popular compiler framework, MLIR (Lattner et al., 2021), to enable seamless mapping of various LLM operations to the PIM command stack. MLIR’s multi-level abstractions, ranging from high-level tensor computations (e.g., linalg.batch_matmul) to hardware-centric operations (e.g., pim.wr_inp), facilitate customized optimizations and transformations for multi-PIM architectures. A runtime launches the compiled PIM command stacks on the multi-PIM environment. We extend IREE (The IREE Authors, 2019) and its hardware abstraction layer (HAL) to integrate PIM’s software development kit (SDK) for memory allocation and command dispatching in end-to-end LLM inference [^1] We implement our LoL-PIM framework using MLIR and IREE Ver. Feb. 18, 2023. (commit eb14186) and (commit 2f40854), respectively. We integrate the AiMX vendor library (AiM-SDK) into our runtime and validate our PIM simulation setup by using AiM-SDK’s software simulator..
Compiling Transformer Variants. Fig. 8 illustrates LoL-PIM’s compilation workflow. The Lowering Pass starts with the processing of Transformer decoder operations represented in the StableHLO dialect (StableHLO, 2019) and translating them into lower-level MLIR dialects like linalg and arith. Mapping Transformer operations of varying sizes and structures is a key challenge. To address this, we implement a Pattern-Matching Pass that extends existing IR traversal to identify sequences of operation types matching predefined decoder patterns like GQA and SwiGLU. In the case of GQA, for example, we detect mismatched vector dimensions between $Q$ and $K,V$. Upon pattern matching, the compiler generates Execution Table specifying the Transformer operation type, partitioning direction, and communication strategy for each operation.
PIM Command Generation. We implemented PIM Dialect in the Codegen Pass to convert the lowered Transformer operations into the PIM command stack. PIM Dialect consists of two main steps: 1) PIM-aware multi-module partitioning and 2) microcoding for PIM command generation. For instance, given the number of PIM modules and workload information (LLM structure, context length, batch size), $QK^{T}$’s GEMV workload is partitioned and distributed to multiple PIM modules (①). For each module, the loop-based PIM command generation step constructs a command stack for the partitioned $QK^{T}$ computation . Since the same computation patterns repeat across Transformer decoder layers and tokens, a PIM command stack is encoded with DPA for compact representation(②). The necessary metadata for runtime dispatching, especially the virtual memory allocation for the weights and KV-cache are also encoded (③).
LoL-PIM Composition 8 Modules, 16 AiM controllers per Module
Frequency 1GHz
Interface Bandwidth 64GB/s
HUB GPR (SRAM) 512KB
Compute Unit Softmax, Layernorm, EWADD, EWMUL
Dispatcher DPA CMD buffer(96KB), Addr Map Buffer(96KB)
AiM Memory Configuration GDDR6 16Gb/s; x16 organization;
512 GB/s; 2 channels per chip; 16 channels;
16 banks/Channel; Page size = 2KB
Processing Unit (PU) 1 PU per bank; 32 GFLOPS per PU(with 1 GHz)
Global Buffer 2KB Global Buffer per channel
Table 5. LoL-PIM’s hardware configurations.
Figure 9. Throughput evaluation of Qwen1.5-7B on standalone (a) and heterogeneous (b) systems. The throughput (tokens/sec) is measured on QMSum HotpotQA and Musique tasks.
[…]
7.2. Hardware Overhead.
To construct the LoL-PIM hardware, a on-module dispatcher is introduced to the HUB to support dynamic memory management, and a I/O-aware buffer is applied to AiM to overcome the I/O bottleneck. To evaluate the overhead of these two hardware components, we use the PIM architecture from (Yongkee Kwon et al., 2024) as the baseline, with the AiM architecture based on information from (Kwon et al., 2023c). The baseline PIM architecture consists of 2 HUBs and a total of 16 AiMs (Yongkee Kwon et al., 2024). To model our LoL-PIM, the on-module dispatcher and I/O-aware buffer are applied to the HUB and AiM, respectively, in the baseline PIM. Since the most dominant hardware component in this PIM architecture is the AiM, the overhead of the I/O-aware buffer is modeled as a percentage of the hardware area of a single AiM chip, while the overhead of the on-module dispatcher is modeled as a percentage of the area corresponding to eight AiMs.
The area of a single AiM, which consists of 16 PUs and 16 bank cells, is approximately 12.16 mm² (Kwon et al., 2023c). Firstly, for the overhead of the I/O-aware buffer, the 2KB area required for additional Global Buffer is modeled using CACTI, resulting in an area of 0.053 mm², which corresponds to an overhead of approximately 0.4%. Secondly, the on-module dispatcher requires a Va2Pa Table, CMD Buffer, configuration buffer, and a decoding unit. The area for the decoding unit is modeled using Synopsys Design Compiler under SAED 32nm technology, while the buffers are modeled using CACTI, resulting in a total area of 0.48 mm². Since one on-module dispatcher is required per eight AiMs, the total overhead for the entire PIM system is approximately 0.5%.
8. Evaluation
8.1. Evaluation Settings
Evaluation LLMs. We evaluate LoL-PIM’s compiler capability using diverse LLMs specified in Table 1, with 1.8B to 72B parameters, considering context lengths up to 32K tokens, and popular activation types (RELU, SwiGLU) implemented in AiM’s LUT. We also consider three tasks with distinct context length characteristics from a representative long-context LLM benchmarks, LongBench (Bai et al., 2024); QMSum (Summarization), HotpotQA and Musique (Multi-Document QA) (Table 2).
Matrix shape (4096, 8192) (4096, 16384) (8192, 4096) (12288, 12288) Cycle Diff. 0.73% 0.79% 0.89% 0.78%
Table 6. Validation of our Ramulator-based PIM simulator (cycle difference with the AiMX vendor library).
GPU-HBM (A100(Nvidia, 2020)) GPU-GDDR6 PIM
Compute Frequency 1.4GHz 1.4GHz 1.0GHz
Peak FLOPS 312T 312T 66T
Off-chip Memory Type HBM3 GDDR6 GDDR6
Capacity 80GB 64GB 64GB
External BW 3352GB/s 4096GB/s 4096GB/s
Internal BW N/A N/A 65.5TB/s
Table 7. Specification of GPU-HBM, GPU-GDDR, and PIM(=LoL-PIM).
Figure 10. Throughput evaluation of Qwen1.5-72B on standalone (a) and heterogeneous (b) systems. The throughput (tokens/sec) is measured on QMSum HotpotQA and Musique tasks.
[…]
Baselines. We evaluate our system against three baselines: (1) a 16-GPU DGX-A100 system, each GPU equipped with 80GB HBM3 and simulated using a DGX simulator (GPU-HBM), (2) 16 AiMX cards, each comprising 8 devices with one Xilinx Virtex UltraScale+ VU9P FPGA and eight 1GB AiM chips utilizing GDDR6 and QSFP links for inter-card communication (PIM), and (3) a GPU-only system using GDDR6 for GPU evaluation with the matched external bandwidth of the PIM system (GPU-GDDR). These baselines were simulated using validated tools, including the DGX simulator (Jaewan Choi, 2024) and our extended version of Ramulator (Kim et al., 2016), which we validated to ensure accuracy (see Table 6). This setup enables a comprehensive comparison of LLM inference performance across architectures. Detailed specifications for the baselines are provided in Table 7.
LoL-PIM. We model LoL-PIM’s latency using a modified Ramulator implementation, validated against AiM-SDK and based on reference design parameters (Table 6). The simulation includes AiM’s core functionality, AiMX’s multicasting interconnect, and power models derived from DRAM references and CACTI. To evaluate performance comprehensively, we test LoL-PIM with three key techniques: ① scalable PIM architecture, ② dynamic PIM memory management, and ③ I/O-aware buffering. Inspired by recent advances in heterogeneous PIM systems (Heo et al., 2024; Park et al., 2024; Seo et al., 2024), we examine LoL-PIM in both standalone and hybrid GPU+PIM configurations (the hybrid configuration halving PIM capacity). All systems, including the baselines and LoL-PIM (①,②,③), are configured to have a matched external memory bandwidth  [^2] The system does not compromise inference accuracy.. Multi-node communication latency is modeled conservatively at 10GB/s, aligned with AiMX card QSFP bandwidth and matched across systems. Table 5 outlines LoL-PIM’s architectural parameters.
8.2. Performance Comparison
Figure 11. Throughput (tokens/sec) measured on Musique with Qwen-7B. The x-axis indicates the combinations of Tensor Parallel (TP) and Pipeline Parallel (PP) across all nodes. The numbers above each point represent the average batch size.
[…]
Figure 12. Latency breakdown of LLM operations on Qwen1.5-72B for LoL-PIM, comparing two configurations. The left plot represents the result in a standalone and the right plot represents the result in a heterogeneous configuration.
[…]
Scalability. We evaluate LoL-PIM’s scalability across three long-context LLM benchmarks(Table 2), focusing on the Qwen1.5-7B model (Fig. 9(a)). The baseline PIM-only system showed the poorest scalability due to inefficient PIM utilization. While it outperformed the GPU-only system at a smaller capacity of 128GB, its scalability limitations led to the lowest performance at the largest capacity of 1024GB, highlighting the drawbacks of conventional PIM mapping approaches at larger scales. In contrast, LoL-PIM demonstrated significantly better scalability. At 1024GB, it outperformed the baseline GPU-GDDR and PIM-only systems by 3.53$\times$ and 4.74$\times$, respectively, showcasing its superior efficiency. This improvement stems from two key factors: (1) pipeline parallelization, which sustains tensor parallelization and prevents PIM channel utilization drops at larger scales, and (2) larger batch sizes enabled by scaling, which reduce pipeline bubble overhead and enhance throughput  [^3] GPU-GDDR is evaluated to match the external bandwidth with PIM-only systems. A similar trend was observed with the Qwen1.5-72B model (Fig. 9(b)). The baseline PIM-only system again exhibited the lowest scalability and was unable to execute the 72B model on a 128GB system due to its parameter size. However, for capacities between 256GB and 1024GB, LoL-PIM consistently outperformed both the baseline GPU-GDDR and PIM-only systems by 8.54$\times$ and 2.65$\times$, respectively. These results underscore LoL-PIM’s capability to efficiently handle larger models with substantial memory and computational demands. Its sustained high throughput across increasing system capacities highlights its robustness and effectiveness in scaling large-scale workloads.
Standalone vs. Heterogeneous Systems. We further compare the scalability of standalone and the heterogeneous systems. Prior works (Heo et al., 2024; Park et al., 2024; Seo et al., 2024) have suggested promising performance of the heterogeneous xPU+PIM architectures, yet they have not been carefully evaluated in the long-context LLM workloads. In Fig. 9 and Fig. 10, we compare the throughput of the standalone and heterogeneous systems for 7B and 72B models, respectively. Each system was evaluated with capacities ranging from 128GB to 1024GB, showcasing the average tokens per second (tok/sec) achieved while processing the evaluation dataset. Although showing scalability in general, GPU+LoL-PIM achieves inferior throughputs compared to the standalone systems (LoL-PIM). This disatifactory performance is due to the long-context LLM’s memory-bounded characteristics; the higher throughput can be achieved with the more internal memory bandwidth. Yet, the speedup of PIM-only system over GPU+PIM is less than the internal memory bandwidth gap (2×2$\times$2 ×) thanks to GPU+PIM’s efficiency in avoiding underutilization for FC layers. Still, the proposed scalable PIM partitioning, dynamic PIM memory management, and I/O-aware buffering consistently improve the GPU+PIM’s performance, similar to the PIM-only systems.
8.3. Ablation Study
Tensor vs. Pipeline Parallelization. LoL-PIM (①) employs a token-parallel approach to maximize pipeline parallelism. However, increasing the number of pipeline stages introduces pipeline bubbles, which reduce utilization. These overheads diminish with larger batch sizes, necessitating different pipeline parallelism strategies depending on the presence of dynamic PIM memory management. Fig. 11 shows how the optimal balance of tensor and pipeline parallelism varies with and without DPA across system scales. With fixed parallelism strategies, DPA improves performance by up to 1.3$\times$. Additionally, under the same DPA condition, performance differences between parallelism combinations can reach 1.73$\times$. The performance gap between optimal parallelism points with and without dynamic memory management is up to 7%, underscoring the sensitivity of LoL-PIM (①) to dynamic adjustments in batch size and parallelism strategies.
Model Scale #PIM Node PIM LoL-PIM (①②) LoL-PIM (①②③)
Token/sec Util. Token/sec Util. Token/sec Util.
7B 4 1833 15.1 2455 20.2 3668 30.1
14B 5 1309 15.4 1737 20.5 2553 30.1
72B 16 737 12.8 1211 21.1 1740 30.3
Table 8. Throughput (token/sec) and utilization (%) for various LLM scales (Qwen series), measured on Musique.
Model Context QMSum HotpotQA Musique Avg. Mistral-7B 32K 25.55 49.77 28.20 34.51 16K 25.36 49.97 27.67 34.33 8K 24.13 43.63 23.49 30.42 4K 21.51 40.30 19.84 27.22 Qwen1.5-7B 32K 23.92 48.11 26.74 32.92 16K 23.32 46.78 28.24 32.78 8K 21.65 44.29 27.64 31.19 4K 20.61 38.05 19.02 25.89
Table 9. Accuracy evaliation of three LongBench tasks with limited context length for Mistral-7B and Qwen1.5-7B.
Latency Breakdown. Fig. 12 highlights LoL-PIM’s effectiveness compared to baseline GPU and PIM systems. For the 72B model, the large weight parameters and KV cache sizes lead to smaller batch sizes, causing memory-bound issues in FC operations. This is particularly evident in hybrid systems, where GPU batch processing provides limited acceleration, resulting in significant FC overhead. LoL-PIM’s token-parallel approach (①) addresses these inefficiencies by reducing memory overhead in the KV cache, minimizing both Attention and FC bottlenecks. Enhancements with (①②) further increase average batch size, with greater benefits observed in heterogeneous systems. Additional latency reduction is achieved through I/O-aware buffering (①②③). Overall, these combined improvements reduce latency by over 60% for both standalone and heterogeneous systems, demonstrating the effectiveness of the proposed methods.
Utilization Trends in Scaling LLM Size. To demonstrate LoL-PIM’s scalability for LLM inference workloads, we focus on utilization trends across models scaling from 7B to 72B parameters compare with baseline PIM-only systems. The node count is proportionally increased to ensure sufficient memory for larger models. The baseline PIM system exhibits consistently low utilization rates, with a pronounced decline as model scale increases, particularly for the 72B model, where utilization drops to 12.8%. In contrast, LoL-PIM (①②③) shows a significant improvement, nearly doubling the utilization of the baseline PIM system and maintaining a stable utilization rate of around 30%, even for larger models. This sustained utilization across different model sizes (as summarized in Table 8) demonstrates the critical role of ITPP, DPA and I/O-aware buffering in achieving scalable LLM inference.
Impact of Context Length on Accuracy in LLM Inference. To justify developing an efficient long-context LLM serving system, we evaluated two popular long-context LLMs from Table 1 on three LongBench tasks: QMSum, HotpotQA, and Musique (context characteristics in Table 2). We used fixed max context lengths ranging from 4K to 32K, truncating exceeding contexts as per LongBench guidelines (Bai et al., 2024). Table 9 shows that context truncation from 4K to 32K significantly degrades LLM inference accuracy by 13.8-29.6%. While shorter contexts enable faster token generation(Fig.3), this speed-up may not justify the accuracy loss in performance-sensitive tasks requiring precise information retrieval from user contexts. This underscores the value of LoL-PIM, a scalable multi-node LLM serving system based on resource-effective, domain-specific PIMs.
9. Related Work
Multi-Node LLM Serving. As LLMs scale, they require more memory and bandwidth, making parallelization techniques like tensor (Shoeybi et al., 2019) and pipeline parallelization (Narayanan et al., 2021) essential for LLM serving. Advanced serving systems (Ten, 2023; Dee, 2023; Cho et al., 2024) now incorporate continuous batching (Yu et al., 2022) and efficient KV cache management (Kwon et al., 2023d) to improve GPU utilization. However, recent studies reveal GPU underutilization in multi-GPU LLM serving, especially during token generation, leading to approaches leveraging hardware heterogeneity (Patel et al., 2024; He and Zhai, 2024). This paper proposes a multi-node, PIM-only LLM serving system to address GPU underutilization and presents an efficient parallelization strategy for multi-node environments.
DRAM-Based PIM. DRAM-based Processing-in-Memory (PIM) has become a key commercial solution for accelerating computational tasks, especially for Large Language Models (LLMs). This includes general-purpose solutions like UPMEM-PIM (Gómez-Luna et al., 2022; Manjunath et al., 2022; Hyun et al., 2023) and domain-specific accelerators such as HBM-PIM (Lee et al., 2021a; Kwon et al., 2021; Kim et al., 2022, 2023, 2024) and AiM (Lee et al., 2022; Kwon et al., 2023b; He et al., 2020), optimized for matrix multiplication. Recent research addresses bottlenecks in homogeneous PIM environments (Zhou et al., 2022; Kal et al., 2023) and heterogeneous xPU-PIM configurations (Park et al., 2024; Heo et al., 2024; Seo et al., 2024; Li et al., 2024; Wang et al., 2022). The introduction of AiMX (Kwon et al., 2022, 2023a), combining FPGA with AiM chips, marks a significant advancement in computational support and parallel processing for LLMs, reflecting the industry’s move towards more efficient, specialized hardware.
10. Conclusions
We introduce LoL-PIM, a multi-node PIM acceleration system addressing these issues. It features advanced parallelism techniques, direct PIM access commands and on-module dispatcher, and I/O-aware buffering for reducing I/O transfer overhead. Evaluations show LoL-PIM significantly enhances throughput and reduces latency for long-context LLM inference, outperforming multi-GPU and GPU-PIM systems (up to 8.54$\times$ and 16.0$\times$ speedup , respectively), enabling more efficient LLM deployment.